If the homozygous dominant genotype makes up 36% of the population what is the frequency of the dominant allele in that population?
1
Expert's answer
2014-10-21T11:09:29-0400
According to the Hardy-Weinberg Law (p^2+2*p*q+q^2=1) a frequency of the homozygous dominant genotype is p^2 = 36%/100= 0.36, therefore the frequency of the dominant allele in that population is the root of p^2=0.6.